OpenStudy (anonymous):

Can someone show the steps on how to establish this identity: 1- (sin²θ/1+cosθ) = cosθ

OpenStudy (reemii):

\(1-\frac{\sin^2 \theta}{1+\cos \theta} = \frac{1+\cos \theta-\sin^2\theta}{1+\cos\theta}\). So, \[ \frac{1+\cos \theta-\sin^2\theta}{1+\cos\theta} = \cos \theta \] if and only if \[ 1+\cos \theta-\sin^2\theta = \cos\theta + \cos^2\theta \] it's almost done now

OpenStudy (raden):

let's try start from left side